For a random sample of 9 women , the average resting pluse rate is x=76 beats per minute ,& the sample standard deviation is s=5 The standard error of the sample mean is? Please explain madam/ sir ..
Answer by ewatrrr(24785)


Hi

Sample size = 9    sample with mean = 76 b/m  and s = 5

Standard Error = s / √n = 5/3 

“standard error” is an estimate of the standard deviation of the sample mean from the true population mean.  The population or sample standard deviation may be used.  

Wish You the Best in your Studies.